# Retrieval-Augmented Generation and Knowledge-Grounded Reasoning for Faithful Patient Discharge Instructions

### v3

@@ Abstract
Language models (LMs), such as ChatGPT, have the potential to assist clinicians in generating various clinical notes. However, LMs are prone to produce “hallucinations”, i.e., generated content that is not aligned with facts and knowledge. In this paper, we propose the Re3Writer method with retrieval-augmented generation and knowledge-grounded reasoning to enable LMs to generate faithful clinical texts. We demonstrate the effectiveness of our method in generating patient discharge instructions. It requires the LMs to understand the patients’ long clinical documents, i.e., the health records during hospitalization, to generate critical instructional information provided both to carers and to the patient at the time of discharge. The proposed Re3Writer imitates the working patterns of physicians to first retrieve related working experience from historical instructions written by physicians, then reason related medical knowledge. Finally, it refines the retrieved working experience and reasoned medical knowledge to extract useful information, which is used to generate the discharge instructions for previously-unseen patients. Our experiments show that, using our method, the performance of five different LMs can be substantially boosted across all metrics. Meanwhile, we show results from human evaluations to measure the effectiveness in terms of fluency, faithfulness, and comprehensiveness.  [^1] The code is available at https://github.com/AI-in-Health/Patient-Instructions
@@ 1 Introduction
To prove the effectiveness of our Re3Writer, we incorporate it into 5 different language models (LMs): 1) recurrent neural network (RNN)-based LM, 2) attention-based LM, 3) hierarchical RNN-based LM, 4) copy mechanism-based LM, 5) fully-attentive LM, i.e., the Transformer [47] LM. The extensive experiments show that our approach can substantially boost the performance of baselines across all metrics.
@@
2 Related Works
The related works are introduced from two aspects: 1) Natural Language Generation (NLG) and 2) Medical Text Generation.
Natural Language Generation (NLG)   It aims to automatically generate coherent text in natural-language from given corresponding input data, which can be, e.g., text [2], images [49], video [48], and audio [9]. Commonly, approaches use an encoder-decoder framework, where the encoder computes intermediate representations for the input source data and the decoder adopts RNNs or CNNs [16, 10], to generate the target sentences given the intermediate representation. The attention mechanism [31, 47, 52] and the copy mechanism [12] have been proposed to directly provide the decoder with the source information, enabling a more efficient use of the source data. In particular, fully-attentive models, such as the Transformer [47], in which no recurrence is required, have successfully achieved state-of-the-art performance for multiple natural language generation tasks. Most recently, the large language models (LLMs) [54], e.g., ChatGPT [33], have been explored to achieve impressive NLG performance. However, most existing work focuses on the natural text-related data, where research concerning the clinical domain remains relatively under-studied.
Medical Text Generation   Recently, language generation in the medical domain has received growing research interest. For example, Jing et al. [20] proposed to generate radiology reports from chest X-ray images; Lee [25] proposed to generate the clinical notes of emergency department cases from discharge diagnosis codes; Ive et al. [19] proposed to generate mental health records. Meanwhile, some works treated the generation of medical text as a summarization task. For example, Zhang et al. [53] proposed to generate clinical impressions by summarizing radiology reports; Scott et al. [38] proposed to generate the patients’ histories by summarizing their respective health records; Several works [35, 13, 30, 11, 39], including HARVEST [15] - a medical summarization tool, proposed to summarize patients’ key conditions from their clinical health records, such as Hassanpour and Langlotz [13] and Shing et al. [39] which summarized important clinical entities from radiology reports and hospital health records, respectively.
Despite the encouraging performances of these methods in generating various types of medical text, to the best of our knowledge, none of them has attempted to generate the Patient Instructions, which has significant implications for supporting care delivery [7, 36, 42] as noted earlier. We here make the first attempt to generate the Patient Instructions, which is typically longer than existing NLG-related output in the medical domain and which covers more diverse topics than previous medical text generation tasks, such as clinical impressions [53], clinical notes [25], and radiology reports [29].
@@ 4.2 Automatic Evaluation
4.3 Human Evaluation
Metrics   We further conduct human evaluation to verify the effectiveness of our approach Re3Writer in clinical practice. To successfully assist physicians and reduce their workloads, it is important to generate accurate patient instructions (faithfulness, precision), such that the model does not generate instructions that “do not exist” according to doctors. It is also necessary to provide comprehensive true instructions (comprehensiveness, recall), i.e., the model does not leave out the important instructions. For example, given the ground truth PI [Instruction_A, Instruction_B] written by physicians, the PI generated by Model_1 is [Instruction_A, Instruction_B, Instruction_C], and the PI generated by Model_2 is [Instruction_A]. Model_1 is better than Model_2 in terms of comprehensiveness, while is worse than Model_2 in terms of faithfulness. Finally, it is unacceptable to generate repeated or otherwise unreadable instructions (fluency). Therefore, we randomly select 200 samples from our PI dataset. The human evaluation is conducted by two junior annotators (medical students) and a senior annotator (clinician). All three annotators have sufficient medical knowledge. By giving the ground-truth PIs, each junior annotator was asked to independently compare the outputs of our approach and that of the baseline models, in terms of the perceived quality of the outputs - including fluency, comprehensiveness, and faithfulness of outputs compared to the corresponding ground-truth instructions. The senior annotator re-evaluated those cases that junior annotators have difficulties deciding. The annotators were blinded to the model that generated the output instructions.
Results   We select a representative baseline, Seq2Seq, and a competitive baseline, Transformer, to show the human evaluation results (Table 4). As may be seen from the results, our Re3Writer is better than baselines with improved performance in terms of the three aspects, i.e., fluency, comprehensiveness and faithfulness. The human evaluation results show that the instructions generated by our approach are of higher clinical quality than the competitive baselines, which proves the advantage of our approach in clinical practice. In particular, by using our Re3Writer, the winning chances increased by a maximum of $43.5-25.5=18$ points and $55-21=34$ points in terms of the faithfulness metric (precision) and comprehensiveness metric (recall), respectively. At last, we further evaluate how many times physicians would have deemed the generated result as “helpful” vs. “unhelpful” in terms of assisting them in writing a PI. The results are reported in Table 4. As we can see, our approach can generate more accurate PIs than the baselines, improving the usefulness of AI systems in better assisting physicians in clinical decision-makings and reducing their workload.
@@ 4.3 Human Evaluation
4.4 Ablation Study
We conduct a quantitative analysis using the Seq2Seq and the Transformer for the purposes of evaluating the contribution of each proposed component: Retrieve, Reason, and Refine.
@@ Acknowledgments
This work was supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre; an InnoHK Project at the Hong Kong Centre for Cerebro-cardiovascular Health Engineering; and the Pandemic Sciences Institute, University of Oxford, Oxford, UK. Fenglin Liu gratefully acknowledges funding from the Clarendon Fund and the Magdalen Graduate Scholarship.

### v4

# Retrieval-Augmented and Knowledge-Grounded Language Models for Faithful Clinical Medicine

@@ Abstract
Language models (LMs), including large language models (such as ChatGPT), have the potential to assist clinicians in generating various clinical notes. However, LMs are prone to produce “hallucinations”, i.e., generated content that is not aligned with facts and knowledge. In this paper, we propose the Re3Writer method with retrieval-augmented generation and knowledge-grounded reasoning to enable LMs to generate faithful clinical texts. We demonstrate the effectiveness of our method in generating patient discharge instructions. It requires the LMs not to only understand the patients’ long clinical documents, i.e., the health records during hospitalization, but also to generate critical instructional information provided both to carers and to the patient at the time of discharge. The proposed Re3Writer imitates the working patterns of physicians to first retrieve related working experience from historical instructions written by physicians, then reason related medical knowledge. Finally, it refines the retrieved working experience and reasoned medical knowledge to extract useful information, which is used to generate the discharge instructions for previously-unseen patients. Our experiments show that, using our method, the performance of five representative LMs can be substantially boosted across all metrics. Meanwhile, we show results from human evaluations to measure the effectiveness in terms of fluency, faithfulness, and comprehensiveness.  [^1] The code is available at https://github.com/AI-in-Health/Patient-Instructions
@@ 1 Introduction
To prove the effectiveness of our Re3Writer, we incorporate it into 5 different language models (LMs) [36]: 1) recurrent neural network (RNN)-based LM, 2) attention-based LM, 3) hierarchical RNN-based LM, 4) copy mechanism-based LM, 5) fully-attentive LM, i.e., the Transformer [33] LM. The extensive experiments show that our approach can substantially boost the performance of baselines across all metrics.
@@ 3.2 Automatic Evaluation
3.3 Human Evaluation
Metrics   We further conduct human evaluation to verify the effectiveness of our approach Re3Writer in clinical practice. To successfully assist physicians and reduce their workloads, it is important to generate accurate patient instructions (faithfulness, precision), such that the model does not generate instructions that “do not exist” according to doctors. It is also necessary to provide comprehensive true instructions (comprehensiveness, recall), i.e., the model does not leave out the important instructions. For example, given the ground truth PI [Instruction_A, Instruction_B] written by physicians, the PI generated by Model_1 is [Instruction_A, Instruction_B, Instruction_C], and the PI generated by Model_2 is [Instruction_A]. Model_1 is better than Model_2 in terms of comprehensiveness, while is worse than Model_2 in terms of faithfulness. Finally, it is unacceptable to generate repeated or otherwise unreadable instructions (fluency). Therefore, we randomly select 200 samples from our PI dataset.
The human evaluation is conducted by two junior annotators (medical students) and a senior annotator (clinician). All three annotators have sufficient medical knowledge. By giving the ground-truth PIs, each junior annotator was asked to independently compare the outputs of our approach and that of the baseline models, in terms of the perceived quality of the outputs - including fluency, comprehensiveness, and faithfulness of outputs compared to the corresponding ground-truth instructions. The senior annotator re-evaluated those cases that junior annotators have difficulties deciding. The annotators were blinded to the model that generated the output instructions.
Results   We select a representative baseline, Seq2Seq, and a competitive baseline, Transformer, to show the human evaluation results (Table 4). As may be seen from the results, our Re3Writer is better than baselines with improved performance in terms of the three aspects, i.e., fluency, comprehensiveness and faithfulness. The human evaluation results show that the instructions generated by our approach are of higher clinical quality than the competitive baselines, which proves the advantage of our approach in clinical practice. In particular, by using our Re3Writer, the winning chances increased by a maximum of $43.5-25.5=18$ points and $55-21=34$ points in terms of the faithfulness metric (precision) and comprehensiveness metric (recall), respectively. At last, we further evaluate how many times physicians would have deemed the generated result as “helpful” vs. “unhelpful” in terms of assisting them in writing a PI. The results are reported in Table 4.
As we can see, our approach can generate more accurate PIs than the baselines, improving the usefulness of AI systems in better assisting physicians in clinical decision-makings and reducing their workload.
@@ 3.3 Human Evaluation
3.4 Ablation Study
We conduct a quantitative analysis using Seq2Seq and Transformer for the purposes of evaluating the contribution of each proposed component: Retrieve, Reason, and Refine.
@@ Acknowledgments
This work was supported in part by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre; an InnoHK Project at the Hong Kong Centre for Cerebro-cardiovascular Health Engineering; and the Pandemic Sciences Institute, University of Oxford, Oxford, UK.